Privacy and Security Notice

FEL Upgrade Project Weekly Brief –July 22-26, 2002



This week we cooled down the newly installed second cryomodule and the revamped injector cryounit. The linac components for the FEL Upgrade are on schedule for the remaining minor hardware additions that will precede testing of the RF hardware over the next few months. Kudos to the SRF, cryo and installation groups plus our own I&C team for their support of this important milestone. Also this week, through the persistent efforts of the W&M/JLab ion implantation team a test tube in the configuration of the cathode support tube was successfully implanted through two complete test runs without arcing using a new low voltage configuration. A new video capture system has been installed by our I&C team that will be used for automated emittance measurements among other useful tasks.


We spent some time analyzing the calculations that are used to support the ODH (oxygen deficiency hazard) classification of the FEL Facility prior to the cool down of the cryomodules on Wednesday July 25. A Temporary Safety Operating Procedure (TSOP) was put in place to allow the installation and cool down operations to proceed until the final configuration of passive and active controls are in place.

Discussions were held with our Engineering Department to estimate our engineering support needs for the remainder of this fiscal year and the next fiscal year for final installation and commissioning activities, in addition to new design tasks associated with the JTO projects and any potential start-up of 100 kW FEL hardware.

On Thursday and Friday we had visitors from AES, Inc for preparation of our presentations to next week’s Navy review of injector systems for high power FELs which will be held at AES’s

Medford, NY facility on July 31-Aug.1.

WBS 4 (Injector):

Continued to work with W&M on the implantation of a test tube. Two runs were completed successfully at half the nominal operating voltage to minimize the possibilities of arcing. No arcs occurred during the runs and the implanted layer is strongly attached to the surface of the tube despite the lower operating voltage. The actual support tube is expected to be ready for implantation at the beginning of next week. The SF6 tank modification drawings were signed off and submitted to the shop. The rear 6-way cross and cathode shield operating bellows were submitted to the shop for final welding.

The light box can is complete. The machine shop is now working on completing the mirror holders that insert into the can. The installation crew has installed two of the four stands for the drive laser transport line.

Gun HVPS - The conditioning resistor for the Gun was fired again to increase its resistance. The value was changed from 65 M ohm to 186 M ohm. This is very close to the desired value. While a section of the resistor was broken off by the heat treat vendor, it is felt that it can be salvaged for a working spare or initial trial unit by (epoxy) gluing the section back into the resistor. Additional material is being ordered to make several more resistors. The door for the HVPS Tank was hung this week. The mating flanges, o-ring and grove are being cleaned and greased.

WBS 6 (RF):

Quarter HVPS - Preparations were made to start testing the HVPS, but sickness of several key people has postponed the testing until next week.

WBS 8 (Instrumentation):

The existing Drive Laser Pulse Controller (DLPC) has been put back together and is ready for FEL commissioning. The upgraded pulse controller will be brought online in parallel with the old system. The first component of the new DLPC has been implemented in iocfel10. It is an F0080 Timer Card which will be used for synthesizing the FEL's pulsed beam repetition rate and handle the beam synchronization requirements detailed in the design review. The DLPC design review presentation can be viewed by clicking link below and a screen capture of the expert EPICS screen for the DLPC's Rep-Rate Timer card as well as a screen-shot from the oscilloscope showing the results can be viewed by clicking the link below. The design for the C0305 (micropulse frequency generator) carrier board is almost finished.

*DLPC design review->http://laser/techdoc/FEL_Project_Review/revans/DLPC_2002_A.htm

*DLPC screen-shots>

All parts are on order to finish-up the F0151 (VME based BLM cards) project. The F0151 layout is nearing completion and is expected to go out for production next week. Additionally, the software to utilize the new hardware is also well in progress.

Control system debug continued on the beamviewer system for the 3F region. Troubleshot and repaired beamviewer chassis' and control cabling. Video cables for the beamviewers have been pulled into position. Video rack (FL02B09) has been disassembled and reworked, hardware in the rack has been tested to accommodate the upgrade.

The new video capture system is up and running. This new system is based on a PC running Linux and acting as an IOC talking directly to EPICS. The PC and capture card are running under Linux now and the EPICS will come on soon. Pavel Evtushenko will be visiting in September and will be working on a fully incorporating this to an automated emittance measurement system. Additionally the PC104 system for control of the optical cavity is also making progress. Vellums for the Fast Raster boards have been received from EECAD and signed off. The raster boards have been sent to the board fabrication house with a 5 day turn around. Six Canbus interface boards were populated. These will be used for the new communications to the beam viewer control chassis. The Quad magnet wiring harness's have been completed for the machine.

WBS 9 (Transport):


Optical Chicane Dipoles (GW)

• At Magnet Enterprises International (MEI) in Oakland CA. The next three GWs are being prepared for shipment the middle of next week.

• The remaining 6 cores of the second batch of 6 Magnets for UV will be ready for gap checking the first week of August.

• The 18th of 20 coils is in potting today.

Injector Dipoles (DU/DV)

• Wang NMR is concentrating on gluing the shim materials to the two GU cores.

• The electrical connectors were received from a local supplier and assembly of GVs was proceeding. They discovered they are missing two of the 24 field clamp parts in the package from Master Machine. We are investigating.

Arc 180 Degree Dipoles (GY)

• At Wang NMR, the fourth GY Coil is being wound.

• GY Coil Potting fixture remains in fabrication.

• Wang has been doing quality checks in anticipation of potting the first three coils. The first coil has not been able to be made short free because of some built-in twist to the conductors formed during the winding process. They believe the coil is not reparable in a reliable way and should be scrapped. The remainder of the coils test out good. They have enough conductor to wind the first two of the four layers of a fifth coil. We ordered more conductor this week to complete the coil.

Arc Bend, Reverse Bend Dipoles (GQ, GX)

• Process Equipment Co. (PECo) is assembling the remaining three GQ cores for an early August shipping date. The GX cores are on track for delivery two weeks later.

• At Wang NMR, the GX coil winding is almost complete. Potting fixtures remain in preparation for the GQ and GX coils. They received 10 gal. of the alternate (non-welting) epoxy formulation from P D. George Co.


Trim Quad (QT)

• The first unit remains in temperature test. We paid for three more units at Milhous but they were held there in anticipation of the cooling plate fix.

Sextupole (SF)

• We finished checking the sextupole drawings from DULY Research. We were disappointed with the quality of the drawings and requested some rework at their expense. We also made substantial improvements to the intra coil connection design (from discussions with MEI) using wrought fittings and the cover. All the corrections were sent to DULY to start the drawing rework.

Octupole (OT)

• Work is just about to be started after DULY signed the Task Order.

Beam Line and Vacuum

• Master machine continued work on the Arc Vacuum Chambers.

• The X Chamber continues fabrication in the Jlab Shop.

• The drawings of the four chambers for the optical chicane region were signed off for Jlab Shop fabrication. The materials have already been purchased.

• Design of the regions around the wiggler continues.

• We signed off the drawings of the stands for sextupoles and trim quads and sent them out for procurement and fabrication.

WBS 11 (Optics):

Last week's tests with the optical cavity hardware turned up another problem with the rotary feedthrough - ~ 10 degrees of backlash before the shaft turned. Discussions with the manufacturer suggest they know the problem existed but it wasn't published in their literature. What concerns us are the consequences of the feedthrough internals failing and rupturing the bellows, and thus venting the chamber. We are revisiting and considering other vendors to find if a better solution exists.

Problems with the motion control software/hardware continue to be worked. We found that an overnight, ~ 100 deg. C bake of a cleaned and disassembled LVDT resulted in a pressure of 3E-8Torr. This is close to the level of performance we need for the IR machine, so we will continue working with this vendor to reduce the outgassing. Tests of the OCMMS beamline hardware show that the mechanical tolerancing was more than adequate. We are in the process of finding the best compromise between ROC measurement accuracy and beam vignetting on the upstream assembly, caused by the requirement to use a 2" dia BPM, rather than the 3" variety.

Tests of the ultrafast laser system diagnostics continue. We are finding that, even an hour after starting the laser, the power doesn't return to the value it had the night before. This is surprising for a temperature-stabilized, diode-pumped system, so we are working with the vendor to determine why this is happening.